OpenStudy (anonymous):

how do you get from y=Ae^-4x + Be^-6x to differential eqution f''(x) +10f(x) +24y=0

OpenStudy (anonymous):

What do you mean by how do you 'get to it?'

OpenStudy (anonymous):

here we have y=f(x) and hence y' = f'(x) and y''=f''(x)

OpenStudy (anonymous):

diiferntiating y we have ( y=Ae^-4x + Be^-6x) y'=-4Ae^-4x -6Be^-6x again on differentiating further we have y''=16Ae^-4x +36Be^-6x now we need to eliminate A and B from y=Ae^-4x + Be^-6x and y'=-4Ae^-4x -6Be^-6x and put the values of A and B in y''=16Ae^-4x +36Be^-6x otherwise y''+24y=16Ae^-4x +36Be^-6x +24Ae^-4x + 24Be^-6x y'' +24y= 10(4Ae^-4x +6Be^-6x) but y'=-4Ae^-4x -6Be^-6x hence y''+24y=-10y' or we have y''+24y+10y'=0 or f''(x)+10f'(x)+24y=0

myininaya (myininaya):

Or since \[y=Ae^{mx}+Be^{nx}\] That means when we solve a quadratic and got \[r=m \text{ or } r=n => r-m=0 \text{ or } r-n=0\] \[=> (r-m)(r-n)=0 => r^2+(-n-m)r+mn=0\] Which means we solve the differential equation: \[y''+(-n-m)y'+mny=0\]